OpenStudy (mikezack123):

Find the exact value of the radical expression in simplest form. the square root of p to the sixth power + the square root of 9 times p to the sixth power + 3 the square root of p to the sixth power − the square root of p squared

OpenStudy (mikezack123):

\[\sqrt{p^6} + \sqrt{9p^6} + 3\sqrt{p^6} - \sqrt{p^2}\]

OpenStudy (mikezack123):

Thats the expression above

OpenStudy (dumbcow):

To add terms you need the exponent to be the same When the power is an even number inside a radical, you can cut it in half and get rid of the radical \[\sqrt{x^{2n}} = x^n\]

OpenStudy (mikezack123):

what do i do next? from there?

OpenStudy (dumbcow):

add like terms \[\rightarrow p^3 + 3p^3 +3p^3 - p\]

OpenStudy (mikezack123):

\[7p^3 - p\]

OpenStudy (mikezack123):

is what i got

OpenStudy (mikezack123):

@dumbcow is that it?